Pompeli, Inc, has sales of $46,500, costs of $21,600, depreciation expense of $1.90o, and interest expense of $1,650. If the tax rate is 21 percent, what is the operating cash flow, or OCF? (Do not round intermediate calculations.) Operating cash flow < Prev 80f 10 İİİ Next > e n e to search
Answer #1

Profit before interest and tax = Sales - Costs - Depreciation

=46500-21600-1900

=23000

Profit before tax = 23000 - 1650

=21350

Taxes = PBT(t)

=21350(0.21)

=4483.5

OCF = EBIT - Taxes + depreciation

=23000 - 4483.5 + 1900

=20,416.5